# Safety-Critical Edge Robotics Architecture with Bounded End-to-End Latency

### v1

@@ II Related Work
inline, color=green!40, author=GG]Estimated size: 0.75 column
@@
inline, color=red!40]GG Abeni and Faggioli [8] investigated the CPU latency of Xen and KVM hypervisors with and without PREEMPT_RT patch for use with RT-cloud nodes. They provided some guidelines for correctly configuring the VMs to reduce the introduced latencies. Several works, such as [9], focused on (hierarchical) container scheduling to support real-time cloud computing. [10] and [7] presented Kubernetes orchestrator extensions to support the deployment of real-time applications (containers) in cloud infrastructure. Gala et al. [1] evaluated existing cloud virtualization technologies for deploying a Real-Time (RT)-cloud to host an existing real-time safety-critical use case. They proposed a resource management layer to use with existing cloud virtualization technology (KVM) to support predictable safety-critical operation as a cloud-based service. Abeni et al. [11] presented the Fault-Tolerant Real-Time Cloud infrastructures capable of hosting highly reliable and real-time applications. Szalay et al. [12] presented an architecture for a real-time Function-as-a-Service platform with the requirements for the underlying network and nodes. The FORA project [13] addressed Industry 4.0 challenges from several angles, such as resource management and middleware, safety and security, and industrial control. inline, color=green!40, author=GG]papers I cited here are mainly aimed at the TPC of the workshop. We will need to expand it for a later conference submission
@@
inline, color=red!40]Johannes, Til There is ongoing research interest in combining industrial robots with cloud/edge computing infrastructure, with several surveys on the topic[14, 3]. Balogh et al. [15] proposed decoupling the closed-loop control of the robot from the robot’s embedded system and placing it into an edge cloud execution environment to benefit from ease of maintenance and improved resiliency to software/hardware failures while allowing the physical platform and control intelligence to evolve separately from each other. Several works (e.g., [4]) have looked into cloud/edge-based robot localization, mapping, or navigation. However, the focus of these works is on performance measures or feasibility, not on the hard real-time constraints on the robot itself. Vick et al. [5] proposed splitting the robotics applications into modular components capable of running on cloud servers while managing hard real-time tasks on the robot, leading to challenges that include guaranteeing reliable low-latency wireless communication[16, 17]. These challenges are in the scope of our proposed approach. Our aim is to dive deeper into the challenges regarding real-time constraints.
inline, color=green!40, author=TU]Vick et al. did a lot in this area, but we should look for other authors, specifically more papers which are not from Telekom or Fraunhofer (or industry). TODO find good source for security related topics
Our related work survey showed differences in computing power installed on the robot. In case of high computing power, the robot may be categorized as an edge device, allowing it to perform complex calculations [18]. In our work, we prioritize the hard real-time requirements and the SWaP constraints of the robot. This means it will use an energy-efficient rather than high-performance architecture, and tasks of lower criticality will be offloaded to a separate edge device. In [18], the authors also utilize the cloud to facilitate collaborative robotics, which is out of scope in our approach. However, we leave it as an option for future work. inline, color=green!40, author=TU]Maybe a different source or more? Extending the previous work on cloud/edge robotics, we tackle this topic with a main focus on the bounded E2E latency, security, and safety aspects.
@@ III Existing Industrial Robotics Use Case
inline, color=red!40]Til, Johannes inline, color=green!40, author=GG]Estimated Size: 1 page Our baseline is a case study from the realm of embedded real-time systems. It consists of a wheeled robot with sensors, a microcontroller with driver software for the hardware components, and a top-level application, which can include any number of extensions, e.g. additional sensors.
@@
inline, color=green!40, author=TU]@GG Feel free to delete the page marker, if it’s not required. We’re citing a fairly large book covering various robotics problems and algorithms, so I tried to be specific.
@@ IV RT-Cloud for Industrial Robotics Use Case
inline, color=green!40, author=GG]Estimated Size: 2 pages inline, color=red!40]GG, LM, IA, IK
@@ IV-A Why use edge-cloud?
inline, color=red!40]GG, TU, IA The particle filter demands high processing power. The main concern is the number of particles. In a complex or highly dynamic environment, one must assume the number of particles to be quite large, and in every step, every particle needs to be traversed by the algorithm. Worse yet, the algorithm must calculate the sensor model’s results for every particle.
@@ IV-C What are the requirements for running in the edge?
inline, color=red!40]GG, Til, LM, IA Robotics applications require dependable and predictable operation with bounded E2E latency. These requirements translate into dealing with faults (e.g., untimely arrival or omission of communicated data), encapsulating applications for safety and security, and providing E2E guarantees.
@@
inline, color=green!40, author=GG]Table with timing requirements/parameters for the SC robotics apps that will run in the edge
@@ IV-D Edge component selection
inline, color=red!40]GG, IK inline, color=green!40, author=GG]add Graph’s from Isser here to motivate the next section Containers have become the de facto industrial approach, especially in cloud environments, providing an ideal underlying layer for edge-to-cloud and multi-cloud scenarios. Hence, we use Linux with Docker containers to separate, isolate, and abstract the safety-critical applications from the COTS hardware. To fully take advantage of the edge node resource and virtualization, we must run multiple robotics and best-effort application containers on the same edge node. However, multiple containers on the same edge node share the underlying resource, such as the network interconnection, last-level cache (LLC), and memory bandwidth. Such resource sharing can cause unpredictable delays and, thus, unbounded E2E latency for robotics applications. Linux and Docker alone are not adequate to meet the requirements for running the robotics application on the edge node and provide bounded E2E latency.
@@
inline, color=red!40]GG The new RMO layer and Linux and Kubernetes extensions are inspired by previous works such as [6, 7]. As suggested in [7], we envisage a RMO layer for the edge that can be integrated with Linux, Docker, and Kubernetes orchestrator. In our RMO layer, each node contains a node-level Local Resource Manager (LRM) to manage its resources in coordination with Linux and Kubernetes. The LRM is comprised of:
@@
inline, color=red!40]LM IEEE 802.11 is the standard that defines the set of protocols necessary for implementing a wireless local area network (WLAN). In a WLAN, nodes communicate with each other over a shared communication medium, where interference and overlapping transmissions can occur, causing frames to have unbounded E2E latencies. In the IEEE 802.11 standard, DCF (Distributed Coordination Function) is the primary method for coordinating node access to the communication medium. DCF employs CSMA-CA (Carrier Sense Multiple Access with Collision Avoidance), which means that a frame transmission only starts after DCF senses that the communication medium is idle. In order to avoid transmission interference, consecutive frames from a node are spaced by bounded delays known as IFS (Interframe Space delays). However, for a time-sensitive application like ours, the collision avoidance capabilities of DCF and the IFS delays are not enough to ensure bounded transmission time for the frames sent (received) by the robot to (from) the cloud.
@@ V End-to-End Guarantees
inline, color=red!40]LM In order to achieve bounded E2E latencies in a WLAN, it is essential that all the possible sources of unbounded delays have been removed from the network or bounded. In the following subsections, we describe which mechanisms need to be added to our WLAN and which modifications in the IEEE 802.11n protocol are needed to upper-bound the E2E latencies.
@@ VI Fault-tolerance and Security Aspects
inline, color=green!40, author=IA]move paragraph below to future work if it sounds a lot We plan to fortify our system against persistent cyberattacks that can successfully compromise the system over time by exhausting the majority of containers in replica groups. This can be achieved by having the RMO layer coordinate the rejuvenation of containers such that any adversarial presence is removed [27]. Furthermore, eliminating adversarial knowledge from previous infiltration attempts requires implementing techniques such as obfuscation [28] and address space randomization [29].
@@ VII Conclusion
inline, color=red!40]GG inline, color=green!40, author=GG]Estimated size: 0.25 column for conclusion inline, color=red!40]GG, IK, LM, IA inline, color=green!40, author=GG]Estimated size: 0.25 column for future work

### v2

@@ II Related Work
Abeni and Faggioli [8] investigated the CPU latency of Xen and KVM hypervisors with and without PREEMPT_RT patch for use with RT-cloud nodes. They provided some guidelines for correctly configuring the VMs to reduce the introduced latencies. Several works, such as [9], focused on (hierarchical) container scheduling to support real-time cloud computing. [10] and [7] presented Kubernetes orchestrator extensions to support the deployment of real-time applications (containers) in cloud infrastructure. Gala et al. [1] evaluated existing cloud virtualization technologies for deploying a Real-Time (RT)-cloud to host an existing real-time safety-critical use case. They proposed a resource management layer to use with existing cloud virtualization technology (KVM) to support predictable safety-critical operation as a cloud-based service. Abeni et al. [11] presented the Fault-Tolerant Real-Time Cloud infrastructures capable of hosting highly reliable and real-time applications. Szalay et al. [12] presented an architecture for a real-time Function-as-a-Service platform with the requirements for the underlying network and nodes. The FORA project [13] addressed Industry 4.0 challenges from several angles, such as resource management and middleware, safety and security, and industrial control.
@@
There is ongoing research interest in combining industrial robots with cloud/edge computing infrastructure, with several surveys on the topic[14, 3]. Balogh et al. [15] proposed decoupling the closed-loop control of the robot from the robot’s embedded system and placing it into an edge cloud execution environment to benefit from ease of maintenance and improved resiliency to software/hardware failures while allowing the physical platform and control intelligence to evolve separately from each other. Several works (e.g., [4]) have looked into cloud/edge-based robot localization, mapping, or navigation. However, the focus of these works is on performance measures or feasibility, not on the hard real-time constraints on the robot itself. Vick et al. [5] proposed splitting the robotics applications into modular components capable of running on cloud servers while managing hard real-time tasks on the robot, leading to challenges that include guaranteeing reliable low-latency wireless communication[16, 17]. These challenges are in the scope of our proposed approach. Our aim is to dive deeper into the challenges regarding real-time constraints.
Our related work survey showed differences in computing power installed on the robot. In case of high computing power, the robot may be categorized as an edge device, allowing it to perform complex calculations [18]. In our work, we prioritize the hard real-time requirements and the SWaP constraints of the robot. This means it will use an energy-efficient rather than high-performance architecture, and tasks of lower criticality will be offloaded to a separate edge device. In [18], the authors also utilize the cloud to facilitate collaborative robotics, which is out of scope in our approach. However, we leave it as an option for future work. Extending the previous work on cloud/edge robotics, we tackle this topic with a main focus on the bounded E2E latency, security, and safety aspects.
@@ III Existing Industrial Robotics Use Case
Our baseline is a case study from the realm of embedded real-time systems. It consists of a wheeled robot with sensors, a microcontroller with driver software for the hardware components, and a top-level application, which can include any number of extensions, e.g. additional sensors.
@@ IV-A Why use edge-cloud?
The particle filter demands high processing power. The main concern is the number of particles. In a complex or highly dynamic environment, one must assume the number of particles to be quite large, and in every step, every particle needs to be traversed by the algorithm. Worse yet, the algorithm must calculate the sensor model’s results for every particle.
@@ IV-C What are the requirements for running in the edge?
Robotics applications require dependable and predictable operation with bounded E2E latency. These requirements translate into dealing with faults (e.g., untimely arrival or omission of communicated data), encapsulating applications for safety and security, and providing E2E guarantees.
@@ IV-D Edge component selection
Containers have become the de facto industrial approach, especially in cloud environments, providing an ideal underlying layer for edge-to-cloud and multi-cloud scenarios. Hence, we use Linux with Docker containers to separate, isolate, and abstract the safety-critical applications from the COTS hardware. To fully take advantage of the edge node resource and virtualization, we must run multiple robotics and best-effort application containers on the same edge node. However, multiple containers on the same edge node share the underlying resource, such as the network interconnection, last-level cache (LLC), and memory bandwidth. Such resource sharing can cause unpredictable delays and, thus, unbounded E2E latency for robotics applications. Linux and Docker alone are not adequate to meet the requirements for running the robotics application on the edge node and provide bounded E2E latency.
@@
The new RMO layer and Linux and Kubernetes extensions are inspired by previous works such as [6, 7]. As suggested in [7], we envisage a RMO layer for the edge that can be integrated with Linux, Docker, and Kubernetes orchestrator. In our RMO layer, each node contains a node-level Local Resource Manager (LRM) to manage its resources in coordination with Linux and Kubernetes. The LRM is comprised of:
@@
IEEE 802.11 is the standard that defines the set of protocols necessary for implementing a wireless local area network (WLAN). In a WLAN, nodes communicate with each other over a shared communication medium, where interference and overlapping transmissions can occur, causing frames to have unbounded E2E latencies. In the IEEE 802.11 standard, DCF (Distributed Coordination Function) is the primary method for coordinating node access to the communication medium. DCF employs CSMA-CA (Carrier Sense Multiple Access with Collision Avoidance), which means that a frame transmission only starts after DCF senses that the communication medium is idle. In order to avoid transmission interference, consecutive frames from a node are spaced by bounded delays known as IFS (Interframe Space delays). However, for a time-sensitive application like ours, the collision avoidance capabilities of DCF and the IFS delays are not enough to ensure bounded transmission time for the frames sent (received) by the robot to (from) the cloud.
@@ V End-to-End Guarantees
In order to achieve bounded E2E latencies in a WLAN, it is essential that all the possible sources of unbounded delays have been removed from the network or bounded. In the following subsections, we describe which mechanisms need to be added to our WLAN and which modifications in the IEEE 802.11n protocol are needed to upper-bound the E2E latencies.
@@ VI Fault-tolerance and Security Aspects
We plan to fortify our system against persistent cyberattacks that can successfully compromise the system over time by exhausting the majority of containers in replica groups. This can be achieved by having the RMO layer coordinate the rejuvenation of containers such that any adversarial presence is removed [27]. Furthermore, eliminating adversarial knowledge from previous infiltration attempts requires implementing techniques such as obfuscation [28] and address space randomization [29].